# Lifshitz transitions and Weyl semimetals from a topological superconductor with supercurrent flow

### v1

@@ Abstract
Similarly, TSs are materials with a huge potential in applications, as they combine two remarkable properties. On the one hand, they host edge modes, known as Majorana quasi-particle (MQPs) [21, 22, 23, 24, 25], featuring peculiar non-local correlations and unconventional braiding properties that could be harnessed for topologically protected quantum computation. On the other hand, they exhibit a dissipationless transport that is ideal to develop green nanoelectronics. Various implementations of TSs, based e.g. on proximized spin-orbit nanowires[26, 27], quantum spin Hall edge systems contacted to ferromagnets[28, 29] and ferromagnetic atom chains[30, 31, 32, 33, 34], have been proposed and are supported by promising, although not yet conclusive, experimental confirmations.[35, 36, 37, 38, 39, 40, 41, 42]
@@
The effect of a superconducting phase modulation is typically neglected in most models of Josephson junctions (JJs), with the argument that current conservation enables one to evaluate the current in the normal weak link, whose properties ultimately determine the critical current. This assumption has also been made in models for JJs[51, 52] or SQUIDs based on $p$-wave TSs.[53, 54, 55] However, because the phase modulation induced by the current modifies the bulk spectrum of the superconductor, the parameter range determining the topological phases of a TS is actually affected by the supercurrent flow itself[56]. In particular, in the last two years the topological effects of a superconducting phase modulation has been investigated for the Kitaev model[57], focusing on the regime where the magnitude $\Delta_{0}$ of superconducting order parameter is larger than the bare bandwidth parameter $w$ [58, 59]. Such a regime $\Delta_{0}>w$, however, is hardly achievable in realistic implementations, and the analysis in more realistic regimes is lacking.
Our analysis overcomes this limitation. Indeed we investigate the effects of the superconducting phase modulation $Q$ on a $p$-wave one dimensional superconductor connected to reservoirs, as sketched in Fig.1. The system is modelled by a Kitaev chain, characterized by a hopping strength $w$, a chemical potential $\mu$, and a magnitude $\Delta_{0}$ of the superconducting order parameter. By analyzing arbitrary parameter values, we show that, while in the regime $\Delta_{0}>w$ the system remains always gapped and the role of $Q$ is to merely modify the values of $\mu$ separating the trivial from the topological phase, a richer scenario emerges for the physically relevant range $\Delta_{0}<w$. In particular, for $|\mu|<2\sqrt{w^{2}-\Delta_{0}^{2}}$ the phase modulation $Q$ can lead the system to become a gapless $p$-wave superconductor. By computing the current in the Kitaev model as a function of the superconducting phase modulation $Q$, we show also that, because of the $p$-wave nature of the superconducting order parameter, the critical value $Q^{*}$ determining the onset of the gapless phase is independent of the chemical potential and only depends on the superconducting order parameter. In our analysis, we determine an upper bound on current that can be driven through the system before the topological Majorana modes disappear.
@@ II Model and symmetries
Model. We consider a one-dimensional $p$-wave superconductor modelled as an infinitely long Kitaev chain, whose Hamiltonian reads
@@
Here $c_{j}\,(c^{\dagger}_{j})$ corresponds to the annihilation (creation) operator, $\mu$ is the chemical potential, and $w>0$ and $\Delta_{0}>0$ are the magnitudes of the hopping amplitude and of the superconducting order parameter, respectively. The wavevector $Q$ (in units of the inverse lattice spacing) characterizing the spatial modulation of the order parameter, describes a net momentum $-2Q$ of the Cooper pair and accounts for the presence of a current flowing through the system.
@@ III Gapped and gapless $p$-wave superconducting phases
showing that the presence of $Q$ makes the two bands no longer symmetric for $k\rightarrow-k$. In Fig.2, the two bands are shown as a function of $k$, at fixed $\mu$ and $\Delta_{0}$ values, for three different values of $Q$. Panel (a) describes the customary $Q=0$ Kitaev model: The two bands are symmetric in $k$, the upper (lower) band is positive (negative) $\forall k$, and a finite direct gap exists between the two bands. As one can see from panel (b) and (c), a finite $Q$ breaks the inversion symmetry [see Eq.(10)] and, when sufficiently large, it can also lead the upper band to acquire negative values, or equivalently, the lower band $E_{-}$ to become positive for the opposite values $-k$. When this occurs, the gap closes indirectly.
Figure 2: The spectrum of the upper band $E_{+}$ (blue dashed curve) and the lower band $E_{-}$ (red curve) for $\Delta_{0}=0.6w$, $\mu=0$. Panels (a), (b) and (c) refer to $Q=0$, $Q=0.1\pi$ and $Q=0.3\pi$.
@@ III.1 The supercurrent carrying state
$\begin{array}[]{lcl}i)&|\mu|>2w\mbox{ \& }&\forall\Delta_{0}>0\mbox{ \& }\forall Q\\ ii)&|\mu|<2w\mbox{ \&}&\sqrt{w^{2}-\mu^{2}/4}<\Delta_{0}\,\mbox{ \& }\cos{Q}\neq|\mu|/2w\\ iii)&|\mu|<2w\mbox{ \& }&w|\sin{Q}|<\Delta_{0}<\sqrt{w^{2}-\mu^{2}/4}\end{array}$ (26)
@@
Gapless superconductor regime. This regime occurs when $E_{+}(k;Q)<0$ (or equivalently $E_{-}(-k;Q)>0$) for some values of $k$, i.e. when the $S_{-}$ sector in Eq.(21) is not empty. Recalling the expression (16), the condition for such gapless superconducting regime to occur are determined by imposing that the condition $|h_{0}(k;Q)|>h(k;Q)$ is fulfilled for some $k$’s. Some lengthy but straightforward algebra [see Appendix A] leads to conclude that such a situation occurs if and only if the following parameter conditions are fulfilled
@@
while the anomalous correlators are
@@ IV Effective 2D fermion model and Lifshitz transition
We now want to discuss how the the presence of the superconducting phase modulation $Q$ affects the topological aspects of the Kitaev model.
@@
The situation is different in the gapless regime, where $E_{+}(k;Q)<0$, for some $k$’s. For such $k$-states the presence of the $h_{0}$-term in Eq.(16) alters the occupancy. In the ground state |$\ket{G(Q)}$ not all the lower band states are occupied, and the approach adopted to topologically classify gapped phases cannot be straightforwardly applied to gapless phases.
However, the presence of the superconducting order parameter modulation $Q$ offers the opportunity to adopt a different perspective. The idea is that $Q$ can be regarded as the wavevector an extra synthetic dimension, in addition to $k$. In this way a 1D superconductor with superconducting phase modulation can be associated to a 2D fermionic model. The Lifshitz transitions[44] occurring in the topology of the Fermi surface $\mathcal{F}$ of the half-filled 2D model enable one to characterize the ground state phases of the 1D superconductor. In particular, we shall show that the $p$-wave superconductor is associated to a 2D WSM, which can be in type-I or type-II regime.
@@ IV.2 Lifshitz transition
Figure 3: The parameter range $2\sqrt{w^{2}-\Delta_{0}^{2}}<|\mu|<2w$. (a) The Fermi surface of the 2D fermionic model associated to the Kitaev chain, with superconducting phase modulation $Q$ treated as a synthetic dimension. The Fermi surface consists of four Weyl nodes, highlighted with black bullets, located at positions given by Eq.(37). Here the plot is given for $\mu>0$. For $\mu<0$ the positions of $Q_{W}$ and $\pi-Q_{W}$ exchange. The green and grey areas denote the gapped topological and trivial phases of the 1D Kitaev chain. The value of $Q_{W}$ determining the topological/trivial boundaries and the $Q$-coordinate of the Weyl nodes is given by Eq.(38) and depends only on on the chemical potential $\mu$ and not on $\Delta_{0}$. (b) Energy band in the vicinity of one of the Weyl nodes, showing that the associated 2D fermionic model is a Type-I WSM.
@@
the Fermi surface consists of four isolated Weyl nodes $\bm{k}_{W}$, where the two bands $E_{+}$ and $E_{-}$ touch, while $E_{+}>0$ and $E_{-}<0$ otherwise ($\bm{k}\neq\bm{k}_{W})$. In this regime the $h_{0}$-term in the Hamiltonian modifies the spectrum but not the occupancy of the bands, and the model (33) corresponds to a two-dimensional Type-I WSM with a Fermi level at $E=0$. The Weyl nodes, shown as black bullets in Fig.3(a) and the type-I band dispersion relation is shown in Fig.3(b).
We shall now argue that this Fermi surface of the 2D model enables us to recover the topological classification of the Kitaev model and to determine the region of parameters where Majorana fermions occur, in the presence of the superconducting phase modulation. Indeed in this regime the lower band $E_{-}$ of the Hamiltonian ${H}$ is completely filled, while the upper band $E_{+}$ is completely empty. Thus, for each cut of the band structure at a fixed $Q$ away from the Weyl nodes ($Q\neq\pm Q^{0,\pi,\pm}_{W}$), one obtains a gapped one-dimensional insulator, which can be classified in the topological class D[63], since it only exhibits the particle-hole symmetry (8). The related topological invariant is the Zak-Berry phase of the lower band,
@@
which is quantized in integer multiples of $\pi$ due to the particle-hole symmetry or, equivalently, through[64]
@@
Trivial phases ($\nu=+1$) correspond to the condition $|\mu|>2w|\cos{Q}|$, whereas topological phases ($\nu=-1$) correspond to $|\mu|<2w|\cos{Q}|$ and, for the case $\mu>0$, are depicted as grey and green regions in Fig.3, respectively. Thus, in the regime (42) the range of $Q$-values where Majorana exist is controlled by the chemical potential $\mu$, and is independent of the value of $\Delta_{0}$. The condition (42) includes the case $\Delta_{0}>w$ studied in Ref.[59]. In the case $\Delta_{0}<w$, Eq.(42) reduces to
@@
which is physically most realistic, as the superconducting order parameter is typically much smaller than the bandwidth parameter $w$. The scenario of the Fermi surface is depicted in Fig.4(a). The Weyl nodes are still located in the positions $\bm{k}^{W}_{j,\pm}$ given by Eqs.(37). However, they are now Type-II nodes, as shown in Fig.4(b), since the Fermi surface of the 2D model exhibits pockets of unpaired electrons and holes, which correspond to the regions where the magnitude of the $h_{0}(\bm{k})$-term overcomes the $h(\bm{k})$-term. Note that electron and hole pockets are mutually mirrors of each other for $k\rightarrow-k$, as a straightforward consequence of the relation (17) originating from the particle-hole constraint Eq.(8) of the BdG formalism. The three panels of Fig.2 represent three cuts of band in the 2D BZ shown in Fig.4(a), at different values of $Q$, where one can see the Lifshitz transition from the gapped phase in panels (a)-(b) (empty Fermi surface) to the gapless phase in panel (c) (appearance of electron and hole pockets), where the ground state of the Kitaev model is given by Eq.(29).
The fermion and hole pockets, identified by $E_{+}(\bm{k})=0$ and $E_{-}(\bm{k})=0$, respectively, cross of the Weyl nodes. As one can see from Fig.4(b), in the range $Q^{*}<|Q|<\pi-Q^{*}$, where
@@
the ground state is gapless, and it exhibits Cooper pairs, unpaired fermions and holes. In contrast, in the range $|Q|<Q^{*}$ and $\pi-Q^{*}<|Q|<\pi$, it is gapped and topological. Indeed the spectrum plotted in Fig.4(c) for a finite-size Kitaev chain of $N_{s}=50$ sites, shows the existence of zero-energy Majorana bound states in such $Q$ range of values.
@@
Figure 4: The parameter range $|\mu|<2\sqrt{w^{2}-\Delta_{0}^{2}}$. (a) The Fermi surface of the 2D fermionic model associated to the Kitaev chain with superconducting phase modulation $Q$ treated as a synthetic dimension. The Fermi surface consists of four Weyl nodes (black bullets), and of electron and hole pockets. The green areas denote the gapped topological phases of the 1D Kitaev chain, while in the other the Kitaev model exhibits a gapless superconducting state, where both Cooper pairs and unpaired fermions and holes are present. The value of $Q^{*}$ determining the topological gapped/gapless boundaries only depends on $\Delta_{0}$ [see Eq.(47)]. (b) energy band in the vicinity of one of the Weyl nodes, showing that the associated 2D fermionic model is a Type-II WSM. Solid and dashed lines enclosing the electron and hole pockets span the entire $k$-direction of the Brillouin zone, and belong to the homology class $(1,0)$. (c) spectrum of the Kitaev chain of $N_{s}=50$ sites for $\mu=w$ and $\Delta_{0}=0.3w$. Zero energy Majorana modes can be seen in the range $|Q|<Q^{*}$ and $\pi-Q^{*}<|Q|<\pi$. (d) $\mu$-$Q$ phase diagram of the system showing the gapped regions and the gapless regions.

### v2

@@ Abstract
Similarly, TSs are materials with a huge potential in applications, as they combine two remarkable properties. On the one hand, they host edge modes, known as Majorana quasi-particle (MQPs) [21, 22, 23, 24, 25], featuring peculiar non-local correlations and unconventional braiding properties that could be harnessed for topologically protected quantum computation. On the other hand, they exhibit a dissipationless transport that is ideal to develop green nanoelectronics. Various implementations of 1D TSs, based e.g. on proximized spin-orbit nanowires[26, 27], quantum spin Hall edge systems contacted to ferromagnets[28, 29] and ferromagnetic atom chains[30, 31, 32, 33, 34, 35], have been proposed and are supported by promising, although not yet conclusive, experimental confirmations.[36, 37, 38, 39, 40, 41, 42, 43]
@@
The effect of a superconducting phase modulation is typically neglected in most models of Josephson junctions (JJs), with the argument that current conservation enables one to evaluate the current in the normal weak link, whose properties ultimately determine the critical current. This assumption has also been made in models for JJs[52, 53] or SQUIDs based on $p$-wave TSs.[33, 54, 55] However, because the phase modulation induced by the current modifies the bulk spectrum of the superconductor, the parameter range determining the topological phases of a TS is actually affected by the supercurrent flow itself[56]. In particular, in the last two years the topological effects of a superconducting phase modulation has been investigated for the Kitaev model[57], focusing on the regime where the magnitude $\Delta_{0}$ of superconducting order parameter is larger than the bare bandwidth parameter $w$ [58, 59]. Such a regime $\Delta_{0}>w$, however, is hardly achievable in realistic implementations, and an investigation in more realistic regimes is lacking.
Our analysis overcomes this limitation. Indeed we investigate the effects of the superconducting phase modulation $Q$ on a $p$-wave 1D superconductor connected to reservoirs, as sketched in Fig.1. The system is modelled by a Kitaev chain, characterized by a hopping strength $w$, a chemical potential $\mu$, and a magnitude $\Delta_{0}$ of the superconducting order parameter. By analyzing arbitrary parameter values, we show that, while in the regime $\Delta_{0}>w$ the system remains always gapped and the role of $Q$ is to merely modify the values of $\mu$ separating the trivial from the topological phase, a richer scenario emerges for the physically relevant range $\Delta_{0}<w$. In particular, for $|\mu|<2\sqrt{w^{2}-\Delta_{0}^{2}}$ the phase modulation $Q$ can lead the system to become a gapless $p$-wave superconductor. By computing the current in the Kitaev model as a function of the superconducting phase modulation $Q$, we show also that, because of the $p$-wave nature of the superconducting order parameter, the critical value $Q^{*}$ determining the onset of the gapless phase is independent of the chemical potential and only depends on the superconducting order parameter. In our analysis, we determine an upper bound on current that can be driven through the system before the topological Majorana modes disappear.
@@ II Model and symmetries
Model. We consider a one-dimensional $p$-wave superconductor modelled as a Kitaev chain, whose Hamiltonian reads
@@
Here $c_{j}\,(c^{\dagger}_{j})$ corresponds to the annihilation (creation) operator at the lattice site $j$, $\mu$ is the chemical potential, and $w>0$ and $\Delta_{0}>0$ are the magnitudes of the hopping amplitude and of the superconducting order parameter, respectively. The wavevector $Q$ (in units of the inverse lattice spacing) characterizing the spatial modulation of the order parameter, describes a net momentum $-2Q$ of the Cooper pair and accounts for the presence of a current flowing through the system.
@@ III Gapped and gapless $p$-wave superconducting phases
showing that the presence of $Q$ makes the two bands no longer symmetric for $k\rightarrow-k$. In Fig.2, the two bands are shown as a function of $k$, at fixed $\mu$ and $\Delta_{0}$ values, for three different values of $Q$. Panel (a) describes the customary $Q=0$ Kitaev model: The two bands are symmetric in $k$, the upper (lower) band is positive (negative) $\forall k$, and a finite direct gap exists between the two bands. As one can see from panel (b) and (c), a finite $Q$ breaks the inversion symmetry [see Eq.(10)] and, when sufficiently large, it can also lead the upper band $E_{+}$ to acquire negative values, or equivalently, the lower band $E_{-}$ to become positive for the opposite values $-k$. When this occurs, the gap closes indirectly.
Figure 2: The spectrum of the upper band $E_{+}$ (blue dashed curve) and the lower band $E_{-}$ (red curve) for $\Delta_{0}=0.6w$, $\mu=0$. Panels (a), (b) and (c) refer to $Q=0$, $Q=0.1\pi$ and $Q=\pi/3$.
@@ III.1 The supercurrent carrying state
$\begin{array}[]{lcl}i)&|\mu|>2w\mbox{ \& }&\forall\Delta_{0}>0\mbox{ \& }\forall Q\\ ii)&|\mu|<2w\mbox{ \&}&\sqrt{w^{2}-\mu^{2}/4}<\Delta_{0}\,\mbox{ \& }|\cos{Q}|\neq|\mu|/2w\\ iii)&|\mu|<2w\mbox{ \& }&w|\sin{Q}|<\Delta_{0}<\sqrt{w^{2}-\mu^{2}/4}\end{array}$ (26)
@@
Gapless superconductor regime. This regime occurs when $E_{+}(k;Q)<0$ (or equivalently $E_{-}(-k;Q)>0$) for some values of $k$, i.e. when the $S_{-}$ sector in Eq.(21) is not empty. Recalling the expression (16), the condition for such gapless superconducting regime to occur are determined by imposing that the condition $|h_{0}(k;Q)|>h(k;Q)$ is fulfilled for some $k$’s. Some lengthy but straightforward algebra [see Appendix A] leads to conclude that such a situation occurs if and only if the following parameter conditions are both fulfilled
@@
while the anomalous correlations are
@@ IV Effective 2D fermion model and Lifshitz transition
We now want to discuss how the presence of the superconducting phase modulation $Q$ affects the topological aspects of the Kitaev model.
@@
The situation is different in the gapless regime, where $E_{+}(k;Q)<0$, for some $k$’s. For such $k$-states the presence of the $h_{0}$-term in Eq.(16) alters the occupancy. In the ground state |$\ket{G(Q)}$ not all the lower band states are occupied, and the approach adopted to topologically classify gapped phases cannot be straightforwardly applied to gapless phases [63].
However, the presence of the superconducting order parameter modulation $Q$ offers the opportunity to adopt a different perspective. The idea is that $Q$ can be regarded as the wavevector of an extra synthetic dimension, in addition to $k$. In this way we can associate a 1D superconductor with superconducting phase modulation to a 2D fermionic model. The Lifshitz transitions[45] occurring in the topology of the Fermi surface $\mathcal{F}$ of the half-filled 2D model enable one to characterize the ground state phases of the 1D superconductor. In particular, we shall show that the $p$-wave superconductor is associated to a 2D WSM, which can be in type-I or type-II regime.
@@ IV.2 Lifshitz transition
Figure 3: The parameter range $2\sqrt{w^{2}-\Delta_{0}^{2}}<|\mu|<2w$. (a) The Fermi surface of the 2D fermionic model associated to the Kitaev chain, with superconducting phase modulation $Q$ treated as a synthetic dimension. The Fermi surface consists of four Weyl nodes, highlighted with black bullets, located at positions given by Eq.(37). Here the plot is given for $\mu>0$. For $\mu<0$ the positions of $Q_{W}$ and $\pi-Q_{W}$ exchange. The green and gray areas denote the gapped topological and trivial phases of the 1D Kitaev chain, respectively. The value of $Q_{W}$ determining the topological/trivial boundaries and the $Q$-coordinate of the Weyl nodes is given by Eq.(38) and depends only on the chemical potential $\mu$ and not on $\Delta_{0}$. (b) Energy band in the vicinity of one of the Weyl nodes, showing that the associated 2D fermionic model is a Type-I WSM.
@@
the Fermi surface consists of four isolated Weyl nodes $\bm{k}_{W}$, where the two bands $E_{+}$ and $E_{-}$ touch, while $E_{+}>0$ and $E_{-}<0$ otherwise ($\bm{k}\neq\bm{k}_{W})$. In this regime the $h_{0}$-term in the Hamiltonian modifies the spectrum but not the occupancy of the bands, and the model (33) corresponds to a 2D Type-I WSM with a Fermi level at $E=0$. The Weyl nodes, shown as black bullets in Fig.3(a) and the type-I band dispersion relation is shown in Fig.3(b).
We shall now argue that this Fermi surface of the 2D model enables us to recover the topological classification of the Kitaev model and to determine the region of parameters where MQPs appear, in the presence of the superconducting phase modulation. Indeed in this regime the lower band $E_{-}$ of the Hamiltonian ${H}$ is completely filled, while the upper band $E_{+}$ is completely empty. Thus, for each cut of the band structure at a fixed $Q$ away from the Weyl nodes ($Q\neq\pm Q^{0,\pi,\pm}_{W}$), one obtains a gapped one-dimensional insulator, which can be classified in the topological class D[64], since it only exhibits the particle-hole symmetry (8). The related topological invariant is the Zak-Berry phase of the lower band,
@@
which is quantized in integer multiples of $\pi$ due to the particle-hole symmetry or, equivalently, the index[65]
@@
Trivial phases ($\nu=+1$) correspond to the condition $|\mu|>2w|\cos{Q}|$, whereas topological phases ($\nu=-1$) correspond to $|\mu|<2w|\cos{Q}|$ and, for the case $\mu>0$, are depicted as gray and green regions in Fig.3, respectively. Thus, in the regime (42) the range of $Q$-values where MQPs exist is controlled by the chemical potential $\mu$, and is independent of the value of $\Delta_{0}$. The condition (42) includes the case $\Delta_{0}>w$ studied in Ref.[59]. In the case $\Delta_{0}<w$, Eq.(42) reduces to
@@
which is physically most realistic, as the superconducting order parameter is typically much smaller than the bandwidth parameter $w$. The scenario of the Fermi surface is depicted in Fig.4(a). The Weyl nodes are still located in the positions $\bm{k}^{W}_{j,\pm}$ given by Eqs.(37). However, they are now Type-II nodes, as shown in Fig.4(b), since the Fermi surface of the 2D model exhibits pockets of unpaired electrons and holes, which correspond to the regions where the magnitude of the $h_{0}(\bm{k})$-term overcomes the $h(\bm{k})$-term. Note that electron and hole pockets are mutually mirrors of each other for $k\rightarrow-k$, as a straightforward consequence of the relation (17) originating from the particle-hole constraint Eq.(8) of the BdG formalism. The three panels of Fig.2 represent three cuts in the 2D BZ shown in Fig.4(a), at different values of $Q$, where one can see the Lifshitz transition from the gapped phase in panels (a)-(b) (empty Fermi surface) to the gapless phase in panel (c) (appearance of electron and hole pockets), where the ground state of the Kitaev model is given by Eq.(29).
The fermion and hole pockets, identified by $E_{+}(\bm{k})=0$ and $E_{-}(\bm{k})=0$, respectively, cross at the Weyl nodes. As one can see from Fig.4(b), in the range $Q^{*}<|Q|<\pi-Q^{*}$, where
@@
the ground state is gapless, and it exhibits Cooper pairs, unpaired fermions and holes. In contrast, in the range $|Q|<Q^{*}$ and $\pi-Q^{*}<|Q|<\pi$, it is gapped and topological. Indeed the spectrum plotted in Fig.4(c) for a finite-size Kitaev chain of $N_{s}=50$ sites, shows the existence of zero-energy MQPs in such $Q$ range of values.
@@
Figure 4: The parameter range $|\mu|<2\sqrt{w^{2}-\Delta_{0}^{2}}$. (a) The Fermi surface of the 2D fermionic model associated to the Kitaev chain with superconducting phase modulation $Q$ treated as a synthetic dimension. The Fermi surface consists of four Weyl nodes (black bullets), and of electron and hole pockets. The green areas denote the gapped topological phases of the 1D Kitaev chain, while in the other the Kitaev model exhibits a gapless superconducting state, where both Cooper pairs and unpaired fermions and holes are present. The value of $Q^{*}$ determining the topological gapped/gapless boundaries only depends on $\Delta_{0}$ [see Eq.(47)]. Solid and dashed lines enclosing the electron and hole pockets span the entire $k$-direction of the Brillouin zone, and belong to the homology class $(1,0)$. (b) energy band in the vicinity of one of the Weyl nodes, showing that the associated 2D fermionic model is a Type-II WSM. (c) spectrum of the Kitaev chain of $N_{s}=50$ sites for $\mu=w$ and $\Delta_{0}=0.3w$. Zero energy Majorana modes can be seen in the range $|Q|<Q^{*}$ and $\pi-Q^{*}<|Q|<\pi$. (d) $\mu$-$Q$ phase diagram of the system showing the gapped regions and the gapless regions, where solid and dashed black lines correspond to the band touching points, i.e. to the location of Type-I and Type-II Weyl nodes $Q_{W}\left(\mu\right)$, respectively [see Eq.(38)].